# Fluid Mechanics Unit Italian Aerospace Research Center (CIRA) Capua (CE), Italy

### v1

@@ 3 Specialized means via finite differences
In the previous section, a correspondence has been mentioned between finite difference discretizations of derivatives of products of two variables and numerical fluxes in the case of two-point discretizations. In particular, the divergence and advective forms with such central or unsymmetric difference operators have been associated with the class of two-point fluxes based on arithmetic averages. In special instances, finite difference splitting can also be used to induce a two-point geometric average [45, 6]—for example, this happens when considering $\phi=(\phi^{1/2}\phi^{1/2})$, in which case $\mathcal{F}(\phi^{1/2},\phi^{1/2})_{i+1/2,\text{adv}}=\mathcal{F}^{\pm}(\phi^{1/2},\phi^{1/2})_{i+1/2,\text{adv}}=\sqrt{\phi_{i}\phi_{i+1}}$. However, in many applications, more general fluxes comprised of nonlinear averages are necessary. These fluxes would seem to be outside the possible schemes that can be generated by classical finite difference discretizations, which are based on linear derivative formulas. For example, with respect to designing EC schemes, one needs access to other nonlinear averages such as logarithmic means. Yet it is has not been evident how to induce such expressions from a splitting technique. It seems then natural to inquire if there is a generalized approach within the finite difference framework to obtaining the discrete equations coming from schemes based on nonlinear flux functions.
By considering the fluxes associated with the general split forms for central or biased derivative operators in Eqs. (35)–(37), one observes that the application of the product rule in manipulating the convective term leads to various formulations which, although equivalent on a continuous ground, are associated with different fluxes with various interpolations that are all based on arithmetic averages for the product of two variables. However, the product rule is a special case of the chain rule, which is also not satisfied by discrete operators, and is associated with the composition of arbitrary nonlinear functions other than the simple product of two variables. From this observation, it is argued that to obtain fluxes in which a nonlinear average arises, of the type of the logarithmic mean appearing in the Ranocha’s EC mass flux [25] ($\mathcal{F}_{\rho}=\overline{\rho}^{(\log)}\,\overline{u}\,$ where $\overline{\rho}^{(\log)}\,=\left(\rho_{i+1}-\rho_{i}\right)/\left(\log\rho_{i+1}-\log\rho_{i}\right)$), nonlinear transformations based on the chain rule have to be used to express the variables and/or their products. These transformations give different expressions that are equivalent on a continuous ground but furnish different approximations when discretely treated. In these discrete forms, the peculiarities of the nonlinear transformation are expected to emerge by determining the form of the fluxes through new nonlinear averages.
@@
Although the equality in Eq. (39) is exact on a continuous ground, the discretization of the right-hand side will be associated with numerical fluxes that are different than those associated with the classical divergence form of the left-hand side. Moreover, using split forms and various choices for the approximations of the derivative of $\rho$ furnishes different expressions for the fluxes, as for the split forms described in Eqs. (35)–(37). This type of procedure, in which FV schemes based on nonlinear fluxes are formulated as FD approximations, can be useful because a generalized treatment of the FD formulations can generate new potentially useful discretizations. An example is studied in Section 4 with reference to EC formulations based on the logarithmic mean.
@@
$\phi=\dfrac{\beta}{\beta+1}\dfrac{\mathrm{d}\phi^{\beta+1}}{\mathrm{d}\phi^{\beta}}.$ (40)
The validity of Eq. (40) is easily verified by observing that the function $\phi^{\beta}$ is a strictly monotone function for positive values of $\phi$. This implies that the relation $\phi^{\beta}=g(\phi)$ can be inverted to give $\phi=g^{-1}(\phi^{\beta})$. For this last point, the inverse function theorem gives
$\dfrac{\mathrm{d}g^{-1}}{\mathrm{d}\phi^{\beta}}=\dfrac{1}{\mathrm{d}g/\mathrm{d}\phi}$ (41)
from which one can manipulate Eq. (40) to give
$f_{\overline{\phi}}(\beta)$$\triangleq$ $\dfrac{\beta}{\beta+1}\dfrac{\mathrm{d}\phi^{\beta+1}}{\mathrm{d}\phi^{\beta}}$
@@
$=$ $\dfrac{\beta}{\beta+1}\left[\left(\beta+1\right)\phi^{\beta}\left(\beta\phi^{\beta-1}\right)^{-1}\right]=\dfrac{\beta}{\beta+1}\left[\left(\dfrac{\beta+1}{\beta}\right)\dfrac{\phi^{\beta}}{\phi^{\beta-1}}\right]=\phi$
Eq. (40) represents the type of fundamental nonlinear transformations used herein to obtain the generalized FD schemes corresponding with more general nonlinear fluxes. Further note that when $\beta=1$, then $\phi$ above may take all real values, both negative and positive. An important observation is that a straightforward evaluation of Eq. (40) breaks down for the cases $\beta=0$ and $\beta=-1$. In such instances, we take the respective limit and employ logarithmic substitutions
@@
$\lim_{\beta\rightarrow-1}\dfrac{\beta}{\beta+1}\left[\left(\dfrac{\beta+1}{\beta}\right)\dfrac{\phi^{\beta}}{\phi^{\beta-1}}\right]=\phi^{-1}/\phi^{-2}=\left(-\dfrac{\mathrm{d}\log(1/\phi)}{\mathrm{d}(\phi)}\right)\left(-\dfrac{\mathrm{d}\phi}{\mathrm{d}1/\phi}\right)=\dfrac{\mathrm{d}\log(1/\phi)}{\mathrm{d}(1/\phi)}$
recovering the type of transformation illustrated with reference to the Ranocha’s mass flux.
Equation (40) therefore re-expresses $\phi$ in terms of a new parameterized function comprised of its derivatives, $\phi=f_{\overline{\phi}}(\beta)$. By substituting finite difference approximations, the limiting case of two-point derivatives obtains various discrete mean values:
@@
    $f_{\overline{\phi}}\left(1\right)=\dfrac{1}{2}\dfrac{\mathrm{d}\phi^{2}}{\mathrm{d}\phi}\approx\dfrac{1}{2}\dfrac{\delta\phi^{2}}{\delta\phi}\ \to\ \dfrac{1}{2}\dfrac{\delta_{\text{2pt}}\phi^{2}}{\delta_{\text{2pt}}\phi}=\dfrac{1}{2}\dfrac{\phi_{a}^{2}-\phi_{b}^{2}}{\phi_{a}-\phi_{b}}=\dfrac{\phi_{a}+\phi_{b}}{2}=\overline{\phi}\,$
@@
    $\beta=\dfrac{1}{2}\quad\rightarrow\quad$Heronian mean:
    $f_{\overline{\phi}}\left(1/2\right)=\dfrac{1}{3}\dfrac{\mathrm{d}\sqrt{\phi}^{3}}{\mathrm{d}\sqrt{\phi}}\approx\dfrac{1}{3}\dfrac{\delta\sqrt{\phi}^{3}}{\delta\sqrt{\phi}}\ \to\ \dfrac{1}{3}\dfrac{\delta_{\text{2pt}}\sqrt{\phi}^{3}}{\delta_{\text{2pt}}\sqrt{\phi}}=\dfrac{1}{3}\dfrac{\left(\sqrt{\phi_{a}}-\sqrt{\phi_{b}}\right)^{3}+3\phi_{a}\sqrt{\phi_{b}}-3\phi_{b}\sqrt{\phi_{a}}}{\sqrt{\phi_{a}}-\sqrt{\phi_{b}}}=\dfrac{\left(\sqrt{\phi_{a}}-\sqrt{\phi_{b}}\right)^{2}+3\sqrt{\phi_{a}}\sqrt{\phi_{b}}}{3}=\dfrac{\phi_{a}+\sqrt{\phi_{a}\phi_{b}}+\phi_{b}}{3}=\overline{\phi}^{E}$
@@
    $\beta=0\quad\rightarrow\quad$Logarithmic mean:
    $f_{\overline{\phi}}\left(0\right)=\dfrac{\mathrm{d}\phi}{\mathrm{d}\log\phi}\approx\dfrac{\delta\phi}{\delta\log\phi}\ \to\ \dfrac{\delta_{\text{2pt}}\phi}{\delta_{\text{2pt}}\log\phi}=\dfrac{\phi_{a}-\phi_{b}}{\log\phi_{a}-\log\phi_{b}}=\overline{\phi}^{(\log)}\,$
@@
    $\beta=-\dfrac{1}{2}\quad\rightarrow\quad$Geometric mean:
    $f_{\overline{\phi}}\left(-1/2\right)=-\dfrac{\mathrm{d}\sqrt{\phi}}{\mathrm{d}(1/\sqrt{\phi})}\approx-\dfrac{\delta\sqrt{\phi}}{\delta(1/\sqrt{\phi})}\ \to\ \dfrac{\delta_{\text{2pt}}\sqrt{\phi}}{\delta_{\text{2pt}}(1/\sqrt{\phi})}=\dfrac{\sqrt{\phi_{a}}-\sqrt{\phi_{b}}}{\dfrac{1}{\sqrt{\phi_{a}}}-\dfrac{1}{\sqrt{\phi_{b}}}}=\sqrt{\phi_{a}\phi_{b}}=\overline{\phi}^{G}$
@@
    $\beta=-1\quad\rightarrow\quad$“Harmonic-logarithmic" mean:
    $f_{\overline{\phi}}\left(-1\right)=\dfrac{\mathrm{d}\log(1/\phi)}{\mathrm{d}(1/\phi)}\approx\dfrac{\delta\log(1/\phi)}{\delta(1/\phi)}\ \to\ \dfrac{\delta_{\text{2pt}}\log(1/\phi)}{\delta_{\text{2pt}}(1/\phi)}=\dfrac{\log(1/\phi_{a})-\log(1\phi_{b})}{\dfrac{1}{\phi_{a}}-\dfrac{1}{\phi_{b}}}=\overline{\phi}^{(H\log)}\,$
  6. 6.
@@
    $f_{\overline{\phi}}\left(-2\right)=2\dfrac{\mathrm{d}\phi^{-1}}{\mathrm{d}\phi^{-2}}\approx\dfrac{\delta\phi^{-1}}{\delta\phi^{-2}}\ \to\ \dfrac{\delta_{\text{2pt}}\phi^{-1}}{\delta_{\text{2pt}}\phi^{-2}}=\dfrac{\dfrac{1}{\phi_{a}}-\dfrac{1}{\phi_{b}}}{\dfrac{1}{\phi_{a}^{2}}-\dfrac{1}{\phi_{b}^{2}}}=2\dfrac{\phi_{b}\phi_{a}}{\phi_{b}+\phi_{a}}=\overline{\phi}^{H}$
Note that in the case of two-point means, we have the property that $f_{\overline{\phi}}(\beta_{a})>f_{\overline{\phi}}(\beta_{b})$ when $\beta_{a}>\beta_{b}$ [46]. Also, specialized procedures for dealing with the risk of a singular denominator will furthermore be necessary in the case of the logarithmic-based means (e.g., Ismail/Roe’s procedure [47]).
Equation (40) thus presents a general parameterized family of means based on derivatives [^5] Another well-known family of means based on differences are the Stolarsky means [48, 49], which also coincide with some of the popular averages mentioned above (e.g., arithmetic, geometric, logarithmic).. In [46], the authors show how to express different means in terms of a quotient expressions comprised of integrals of $\phi$. While the present work independently developed the current derivative-based representation, we acknowledge the previous work’s guidance towards the generalization provided in Eq. (40).
@@
As will be shown in the following section, a direct FD treatment of these analytic forms results in discrete entropy conservation for the current calorically-perfect setting with entropy defined in Eq. (6).
@@ 4 A class of finite-difference entropy-conserving methods
Note that the above leverages the energy consistent relations from Eq. (4) and (5). The current intent is then to recover the entropy conservation statement by judiciously discretizing the primary system. To this end, we first assume a direct discretization of the respective convective and pressure terms previously given in Eqs. (1)–(5). The centered and biased FD stencil renditions for each term associated with the $(n)$ direction derivative are given in the following Tables 1 and 2, where the $\bm{\mathcal{C}}$ and $\bm{\mathcal{P}}$ symbols can now refer to discrete representations based on the context. We allow for additional flexibility in how the density and internal energy will be defined via the specialized FD means presented in Section 3; the necessary forms for achieving exact entropy conservation are derived below. Note that the choice of $\bm{\mathcal{C}}_{\rho}$ automatically specifies the momentum treatment $\bm{\mathcal{C}}_{\rho u}$ that is necessary for fulfilling kinetic-energy-preservation. As we presume a divergence like discretization of mass, the associated KEP form is that of Feiereisen [50].
@@
$\bm{\mathcal{C}}_{\rho e}^{(n)}$ $\delta_{x_{n}}^{\text{c}}(\tilde{\rho}u_{n}\tilde{e})$
$\bm{\mathcal{P}}_{\rho u_{m}}^{(n)}$ $\delta_{x_{m}}^{\text{c}}p$ if $n=m$, otherwise 00
@@
Table 1: Finite-difference discretization of Euler derivatives assuming SBP-compatible stencils. In the case of the FD-EC methods with central convective operators, $\bm{\mathcal{C}}^{(n)}$, the auxiliary variables are $\tilde{\rho}\equiv\frac{\delta_{x_{n}}^{\text{c}}\rho}{\delta_{x_{n}}^{\text{c}}\log\rho}\triangleq\overline{\rho}^{(\log,n)}$ and $\tilde{e}\equiv\frac{\delta_{x_{n}}^{\text{c}}\log e^{-1}}{\delta_{x_{n}}^{\text{c}}e^{-1}}\triangleq\overline{e}^{(H\log,n)}$.
@@
$\bm{\mathcal{C}}_{\rho}^{(\pm,n)}$ $\delta^{\pm}_{x_{n}}(\tilde{\rho}u_{n})$
$\bm{\mathcal{C}}_{\rho u_{m}}^{(\pm,n)}$ $\frac{1}{2}\left[\delta_{x_{n}}^{\pm}(\tilde{\rho}u_{n}u_{m})+\tilde{\rho}u_{n}\delta_{x_{n}}^{\mp}u_{m}+u_{m}\delta_{x_{n}}^{\pm}(\tilde{\rho}u_{n})\right]$
$\bm{\mathcal{C}}_{\rho e}^{(\pm,n)}$ $\delta^{\pm}_{x_{m}}(\tilde{\rho}u_{n}\tilde{e})$
$\bm{\mathcal{P}}_{\rho u_{m}}^{(\pm,n)}$ $\delta^{\mp}_{x_{m}}p$ if $n=m$, otherwise 00
$\bm{\mathcal{P}}_{\rho e}^{(\pm,n)}$ $p\delta^{\pm}_{x_{n}}u_{n}$
Table 2: Finite-difference discretization of Euler derivatives assuming SBP-compatible stencils. In the case of the FD-EC methods with biased convective operators, $\bm{\mathcal{C}}^{(\pm,n)}$, the auxiliary variables are $\tilde{\rho}\equiv\frac{\delta_{x_{n}}^{\mp}\rho}{\delta_{x_{n}}^{\mp}\log\rho}\triangleq\overline{\rho}^{(\log,\mp,n)}$ and $\tilde{e}\equiv\frac{\delta^{\mp}\log e^{-1}}{\delta^{\mp}e^{-1}}\triangleq\overline{e}^{(H\log,\mp,n)}$.
Entropy conservation is enforced by choosing $\tilde{\rho}$ and $\tilde{e}$ such that any spurious volumetric terms are canceled and all that remain are consistent boundary terms. This is required in a point-wise manner in order to further guarantee local conservation in the discrete entropy dynamics. To see this, one employs the presumed form of the discretized convective and pressure terms while furthermore employing the calorically perfect gas assumptions with respect to the definition of entropy, as well as re-writing $1/T=c_{v}/e$; these manipulations yield the following (presented in one-dimension and the right-biased velocity rendition for simplicity):

### v2

@@ 3 Specialized means via finite differences
In the previous section, a correspondence has been mentioned between finite difference discretizations of derivatives of products of two variables and numerical fluxes in the case of two-point discretizations. In particular, the divergence and advective forms with such central or non-symmetric difference operators have been associated with the class of two-point fluxes based on arithmetic averages. In special instances, finite difference splitting can also be used to induce a two-point geometric average [56, 6]—for example, this happens when considering $\phi=(\phi^{1/2}\phi^{1/2})$, in which case $\mathcal{F}(\phi^{1/2},\phi^{1/2})_{i+1/2,\text{adv}}=\mathcal{F}^{\pm}(\phi^{1/2},\phi^{1/2})_{i+1/2,\text{adv}}=\sqrt{\phi_{i}\phi_{i+1}}$. However, in many applications, more general fluxes comprised of nonlinear averages are necessary. These fluxes would seem to be outside the possible schemes that can be generated by classical finite difference discretizations. For example, with respect to designing EC schemes, one needs access to other nonlinear averages such as logarithmic means. Yet, it has not been evident how to induce such expressions from a splitting technique as is possible with respect to the geometric mean. It then seems natural to inquire if there is a generalized approach within the finite-difference framework to obtain the discrete equations coming from schemes based on nonlinear flux functions.
From Eqs. (35)–(37), one observes that the application of the product rule for manipulating the convective term leads to various flux formulations that are based on arithmetic averages for the product of two variables. However, the product rule is a special case of the chain rule, which is not satisfied by discrete operators and is associated with the composition of arbitrary nonlinear functions other than the simple product of two variables. From this observation, it is argued that to obtain fluxes in which a nonlinear average arises (e.g., of the type of the logarithmic mean appearing in the Ranocha’s EC mass flux [30, 31]: $\mathcal{F}_{\rho}=\overline{\rho}^{(\log)}\,\overline{u}\,$ where $\overline{\rho}^{(\log)}\,=\left(\rho_{i+1}-\rho_{i}\right)/\left(\log\rho_{i+1}-\log\rho_{i}\right)$), nonlinear transformations based on the chain rule have to be leveraged to re-express the variables and/or their products. These transformations could then give different expressions that are equivalent on a continuous ground but that furnish different approximations when treated discretely. In these discrete forms, the peculiarities of the nonlinear transformation are expected to influence the form of the fluxes and induce new nonlinear averages, as motivated below.
@@
Although the equality in Eq. (39) is exact on a continuous ground, the discretization of the right-hand side will be associated with numerical fluxes that are different than those associated with the classical divergence form of the left-hand side. In fact, a coordinated discretization of the internal and external derivatives in Eq. (39) can be shown to be equivalent to a FV formulation based on the Ranocha’s mass flux (cf. Section 4). Moreover, using split forms and various choices for the approximations of the derivative of $\rho$ furnishes different expressions for the fluxes, as for the split forms described in Eqs. (35)-(37). This type of procedure, in which FV schemes based on nonlinear fluxes are formulated as FD approximations, can be useful because a generalized treatment of the FD formulations can generate new potentially useful discretizations. An example is studied in Section 4 with reference to EC formulations based on the logarithmic mean.
@@
$\phi$ $=$ $\dfrac{\beta}{\beta+1}\dfrac{\mathrm{d}\phi^{\beta+1}}{\mathrm{d}\phi^{\beta}}$
@@
$=$ $\dfrac{\beta}{\beta+1}\left[\left(\beta+1\right)\phi^{\beta}\left(\beta\phi^{\beta-1}\right)^{-1}\right]=\dfrac{\beta}{\beta+1}\left[\left(\dfrac{\beta+1}{\beta}\right)\dfrac{\phi^{\beta}}{\phi^{\beta-1}}\right]=\phi.$
The validity of Eq. (3) is easily verified by observing that the function $\phi^{\beta}$ is a strictly monotone function for positive values of $\phi$. Eq. (3) represents the type of fundamental nonlinear transformations used herein to obtain the generalized FD schemes corresponding with more general nonlinear fluxes. Further note that when $\beta=1$, then $\phi$ above may take all real values, both negative and positive. An important observation is that a straightforward evaluation of Eq. (3) breaks down for the cases $\beta=0$ and $\beta=-1$. In such instances, we take the respective limit and employ logarithmic substitutions
@@
$\lim_{\beta\rightarrow-1}\dfrac{\beta}{\beta+1}\left[\left(\dfrac{\beta+1}{\beta}\right)\dfrac{\phi^{\beta}}{\phi^{\beta-1}}\right]=\phi^{-1}/\phi^{-2}=\left(-\dfrac{\mathrm{d}\log(1/\phi)}{\mathrm{d}(\phi)}\right)\left(-\dfrac{\mathrm{d}\phi}{\mathrm{d}1/\phi}\right)=\dfrac{\mathrm{d}\log(1/\phi)}{\mathrm{d}(1/\phi)}$ (42)
recovering the type of transformation illustrated with reference to the Ranocha’s mass and internal energy flux.
Equation (3) therefore re-expresses $\phi$ in terms of a new parameterized function comprised of its derivatives, $\phi=f_{\overline{\phi}}(\beta)$.
Substituting finite difference approximations yields a consistent discrete filtered quantity
$f_{\overline{\phi}}(\beta)\triangleq\dfrac{\beta}{\beta+1}\dfrac{\delta\phi^{\beta+1}}{\delta\phi^{\beta}}\approx\phi\ ,$ (43)
with the limiting case of two-point derivatives recovering various discrete mean values:
@@
    $f_{\overline{\phi}}\left(1\right)=\dfrac{1}{2}\dfrac{\delta\phi^{2}}{\delta\phi}\ \to\ \dfrac{1}{2}\dfrac{\delta_{\text{2pt}}\phi^{2}}{\delta_{\text{2pt}}\phi}=\dfrac{1}{2}\dfrac{\phi_{a}^{2}-\phi_{b}^{2}}{\phi_{a}-\phi_{b}}=\dfrac{\phi_{a}+\phi_{b}}{2}=\overline{\phi}\,$
@@
    $\beta=0\quad\rightarrow\quadLogarithmic$ mean:
    $f_{\overline{\phi}}\left(0\right)=\dfrac{\delta\phi}{\delta\log\phi}\ \to\ \dfrac{\delta_{\text{2pt}}\phi}{\delta_{\text{2pt}}\log\phi}=\dfrac{\phi_{a}-\phi_{b}}{\log\phi_{a}-\log\phi_{b}}=\overline{\phi}^{(\log)}\,$
@@
    $\beta=-\dfrac{1}{2}\quad\rightarrow\quadGeometric$ mean:
    $f_{\overline{\phi}}\left(-1/2\right)=-\dfrac{\delta\sqrt{\phi}}{\delta(1/\sqrt{\phi})}\ \to\ \dfrac{\delta_{\text{2pt}}\sqrt{\phi}}{\delta_{\text{2pt}}(1/\sqrt{\phi})}=\dfrac{\sqrt{\phi_{a}}-\sqrt{\phi_{b}}}{\dfrac{1}{\sqrt{\phi_{a}}}-\dfrac{1}{\sqrt{\phi_{b}}}}=\sqrt{\phi_{a}\phi_{b}}=\overline{\phi}^{G}$
@@
    $\beta=-1\quad\rightarrow\quad$“Harmonic-logarithmic" mean:
    $f_{\overline{\phi}}\left(-1\right)=\dfrac{\delta\log(1/\phi)}{\delta(1/\phi)}\ \to\ \dfrac{\delta_{\text{2pt}}\log(1/\phi)}{\delta_{\text{2pt}}(1/\phi)}=\dfrac{\log(1/\phi_{a})-\log(1\phi_{b})}{\dfrac{1}{\phi_{a}}-\dfrac{1}{\phi_{b}}}=\overline{\phi}^{(H\log)}\,$
@@
    $f_{\overline{\phi}}\left(-2\right)=\dfrac{\delta\phi^{-1}}{\delta\phi^{-2}}\ \to\ \dfrac{\delta_{\text{2pt}}\phi^{-1}}{\delta_{\text{2pt}}\phi^{-2}}=\dfrac{\dfrac{1}{\phi_{a}}-\dfrac{1}{\phi_{b}}}{\dfrac{1}{\phi_{a}^{2}}-\dfrac{1}{\phi_{b}^{2}}}=2\dfrac{\phi_{b}\phi_{a}}{\phi_{b}+\phi_{a}}=\overline{\phi}^{H}$
Other means may be recovered by the above formulas, such as the the Heronian mean when $\beta=1/2$. Also note that in the case of two-point means, we have the property that $f_{\overline{\phi}}(\beta_{a})>f_{\overline{\phi}}(\beta_{b})$ when $\beta_{a}>\beta_{b}$ [57]. Also, specialized procedures for dealing with the risk of a singular denominator will furthermore be necessary in the case of the logarithmic-based means (e.g., Ismail/Roe’s procedure [58]).
Equation (3) thus presents a general parameterized family of filters based on derivatives, where a mean of the inputs is recovered in the special case of two-point stencils. To see this latter property, the two-point discretization of $f_{\bar{\phi}}(\beta)$ can be re-written in integral form [^5] Another well-known class of means based on differences are the Stolarsky means [59, 60], which also coincide with some of the popular averages mentioned above (e.g., arithmetic, geometric, logarithmic). The Stolarski means induce an integral form, $\overline{\phi^{\beta}}\cdot\int_{1}^{2}\,\mathrm{d}\phi=\int_{1}^{2}\phi^{\beta}\,\mathrm{d}\phi$ but with the additional approximation that $\overline{\phi^{\beta}}\approx\bar{\phi}^{\beta}$. as
$\bar{\phi}\cdot\frac{1}{\beta}\cdot\delta_{\text{2pt}}\phi^{\beta}=\frac{1}{\beta+1}\cdot\delta_{\text{2pt}}\phi^{\beta+1}\quad\to\quad\bar{\phi}\cdot\int_{1}^{2}\phi^{\beta-1}\,\mathrm{d}\phi=\int_{1}^{2}\phi^{\beta}\,\mathrm{d}\phi\ .$ (44)
Thus we see that $\bar{\phi}$ results from the use of the intermediate value theorem and therefore should lie within the interval $\phi_{1},\phi_{2}]$. We note that the above recovers the same two-point means proposed in [57]  [^6] While the present work independently developed the current derivative-based representation, we acknowledge the work of [57] towards the generalization provided in Eq. (3).; however, the differential perspective offered by Eq. (3) directly generalizes to difference stencils of arbitrary length (e.g., high-order differencing), albeit coming at the loss of guaranteeing that the resulting filtered variable is bounded by its inputs.
@@
As will be shown in the following section, a direct FD treatment of these analytic forms results in discrete entropy conservation for the current calorically perfect setting with entropy defined in Eq. (6).
@@ 4 A class of finite-difference entropy-conserving methods
Note that the above leverages the energy consistent relations from Eq. (4) and (5), namely that the discretization of total energy is comprised of an internal energy discretization plus a kinetic energy component that stems directly from the discretized mass and momentum equations. The current intent is then to recover the entropy conservation statement by judiciously discretizing the primary system. To this end, we first assume a direct discretization of the respective convective and pressure terms previously given in Eqs. (1)–(5). The centered and biased FD stencil renditions for each term associated with the $(n)$ direction derivative are given in the following Tables 1 and 2, where the $\bm{\mathcal{C}}$ and $\bm{\mathcal{P}}$ symbols can now refer to discrete representations based on the context. We allow for additional flexibility in how the density and internal energy will be defined via the specialized FD means presented in Section 3; the necessary forms for achieving exact entropy conservation are derived below. Note that the choice of $\bm{\mathcal{C}}_{\rho}$ automatically specifies the momentum treatment $\bm{\mathcal{C}}_{\rho u}$ that is necessary for fulfilling kinetic-energy-preservation (KEP). As we presume a divergence-like discretization of mass, the associated KEP form is that of Feiereisen [61]. Meanwhile the internal energy convective term is also in divergence form and therefore will naturally admit pressure-equilibrium preservation. These scheme properties are further discussed in Section 4.1. Note that while the split forms presented in the tables are written relative to the local stencil, a matrix-vector rendition stems naturally, as utilized in Appendix A (e.g., $a\delta_{x}b\to[\text{diag}\{{\bf a}\}]{{\sf D}}_{x}{\bf b}$). Alternatively, Appendix B provides the methods in a Hadamard product form.
@@
[…]
$\bm{\mathcal{P}}_{\rho u_{m}}^{(n)}$ $\delta_{x_{m}}^{\text{c}}p$ if $n=m$, otherwise 0
@@
Table 1: A kinetic-energy and pressure-equilibrium preserving finite-difference discretization of Euler derivatives assuming SBP-compatible stencils. In the case of the FD-EC methods with central convective operators, $\bm{\mathcal{C}}^{(n)}$, the auxiliary variables are $\tilde{\rho}\equiv\frac{\delta_{x_{n}}^{\text{c}}\rho}{\delta_{x_{n}}^{\text{c}}\log\rho}\triangleq\overline{\rho}^{(\log,n)}$ and $\tilde{e}\equiv\frac{\delta_{x_{n}}^{\text{c}}\log e^{-1}}{\delta_{x_{n}}^{\text{c}}e^{-1}}\triangleq\overline{e}^{(H\log,n)}$.
@@
$\bm{\mathcal{C}}_{\rho}^{(n,\pm)}$ $\delta^{\pm}_{x_{n}}(\tilde{\rho}u_{n})$
$\bm{\mathcal{C}}_{\rho u_{m}}^{(n,\pm)}$ $\frac{1}{2}\left[\delta_{x_{n}}^{\pm}(\tilde{\rho}u_{n}u_{m})+\tilde{\rho}u_{n}\delta_{x_{n}}^{\mp}u_{m}+u_{m}\delta_{x_{n}}^{\pm}(\tilde{\rho}u_{n})\right]$
[…]
$\bm{\mathcal{P}}_{\rho u_{m}}^{(n,\pm)}$ $\delta^{\mp}_{x_{m}}p$ if $n=m$, otherwise 0
$\bm{\mathcal{P}}_{\rho e}^{(n,\pm)}$ $p\delta^{\pm}_{x_{n}}u_{n}$
Table 2: A kinetic-energy and pressure-equilibrium preserving finite-difference discretization of Euler derivatives assuming SBP-compatible stencils. In the case of the FD-EC methods with biased convective operators, $\bm{\mathcal{C}}^{(n,\pm)}$, the auxiliary variables are $\tilde{\rho}\equiv\frac{\delta_{x_{n}}^{\mp}\rho}{\delta_{x_{n}}^{\mp}\log\rho}\triangleq\overline{\rho}^{(\log,n,\mp)}$ and $\tilde{e}\equiv\frac{\delta^{\mp}\log e^{-1}}{\delta^{\mp}e^{-1}}\triangleq\overline{e}^{(H\log,n,\mp)}$.
The additional entropy conservative property is enforced by choosing $\tilde{\rho}$ and $\tilde{e}$ such that any spurious volumetric terms are canceled and all that remain are consistent boundary terms. This is required in a point-wise manner in order to further guarantee local conservation in the discrete entropy dynamics. To see this, one employs the presumed form of the discretized convective and pressure terms while furthermore employing the calorically perfect gas assumptions with respect to the definition of entropy, as well as re-writing $1/T=c_{v}/e$; these manipulations yield the following (presented in one-dimension and the right-biased velocity rendition for simplicity):